# Positive mass and isoperimetry for continuous metrics with nonnegative scalar curvature

### v2

@@ Appendix A Auxiliary results
Boundedness of isoperimetric sets usually follows from a suitable deformation lemma, see e.g. [AFP21, Theorem B.1]. It is not completely clear how to work out such an argument in the $C^{0}$-setting, yet one can slightly modify the argument, taking advantage of the asymptotic behavior imposed, to obtain the following weaker version, that still suffices for our aims.
@@
Without loss of generality, we can assume that $E$ has positive volume. Take $p\in M$. Let $V(r)=|{E\setminus B_{r}(p)}|$. Call $A(r)=P(E,M\setminus B_{r}(p))$. By coarea (see Theorem 2.1), and since by Appendix A we have $|D\mathsf{d}_{p}|=\mathcal{H}^{n}$, notice that $V^{\prime}(r)=-P(B_{r}(p),E)$. Moreover, a Euclidean-like isoperimetric inequality holds for small volumes. Indeed, the proof of Section 2.3 (compare also with item (1) in Section 2.3) shows that there is a constant $\zeta$ such that $|B_{r}(p)|\geq\zeta r^{3}$ for every $r\in(0,1]$, and every $p\in M$. Thus [APP, Proposition 3.20] gives that $I(v)\geq\xi v^{2/3}$ for every $v\in(0,\eta)$ for some $\eta,\xi>0$. Now arguing verbatim as after [AFP21, Equation (B.1)], we can reduce ourselves to show that, for some constant $C$, there holds
@@
for almost all large enough $r$. Assuming by contradiction that $V(r)>0$ for any $r>0$, the latter would result in a contradiction by ODE comparison on $V(r)$, see [AFP21, Theorem B.1]. To this aim, let $B^{r}$ be a smooth deformation of $B\subset\mathbb{R}^{n}$ of volume $\left\lvert B\right\rvert+V(r)$ such that
@@
where $C$ only depends on $B$. Notice it is enough to take $B^{r}$ to be a ball containing $B$. Then, $\left\lvert E\cap B_{r}(p)\right\rvert+\left\lvert B^{r}\right\rvert_{\mathrm{eu}}=V$ and thus, by using the analogue of [AFP21, Proposition 3.2], which can be proved analogously as in [AFP21] (again exploiting that, arguing as in [AFP21, Lemma 2.17], the isoperimetric profile is achieved by bounded sets), we get
@@
On the other hand, for almost every radius, we have (see [APPV23, Proposition 2.6], which can be applied thanks to Section 2.3) that
@@
[…]

### v3

@@ Appendix A Auxiliary results
Boundedness of isoperimetric sets usually follows from a suitable deformation lemma, see e.g. [4, Theorem B.1]. It is not completely clear how to work out such an argument in the $C^{0}$-setting, yet one can slightly modify the argument, taking advantage of the asymptotic behavior imposed, to obtain the following weaker version, that still suffices for our aims.
@@
Without loss of generality, we can assume that $E$ has positive volume. Indeed, if $|E|=0$, then $\emptyset$ is a bounded representative. Take $p\in M$. Let $V(r)=|{E\setminus B_{r}(p)}|$. Call $A(r)=P(E,M\setminus B_{r}(p))$. By coarea (see Theorem˜2.7), and since by Appendix˜A we have $|D\mathsf{d}_{p}|=\mathcal{H}^{n}$, notice that $V^{\prime}(r)=-P(B_{r}(p),E)$. Moreover, a Euclidean-like isoperimetric inequality holds for small volumes. Indeed, the proof of Section˜2.3 (compare also with item (1) in Section˜2.3) shows that there is a constant $\zeta$ such that $|B_{r}(p)|\geq\zeta r^{3}$ for every $r\in(0,1]$, and every $p\in M$. Thus [7, Proposition 3.20] gives that $I(v)\geq\xi v^{2/3}$ for every $v\in(0,\eta)$ for some $\eta,\xi>0$. Now arguing verbatim as after [4, Equation (B.1)], we can reduce ourselves to show that, for some constant $C$, there holds
@@
for almost all large enough $r$. Assuming by contradiction that $V(r)>0$ for any $r>0$, the latter would result in a contradiction by ODE comparison on $V(r)$, see [4, Theorem B.1]. To this aim, let $B^{r}$ be a smooth deformation of $B\subset\mathbb{R}^{n}$ of volume $\left\lvert B\right\rvert+V(r)$ such that
@@
where $C$ only depends on $B$. Notice it is enough to take $B^{r}$ to be a ball containing $B$. Then, $\left\lvert E\cap B_{r}(p)\right\rvert+\left\lvert B^{r}\right\rvert_{\mathrm{eu}}=V$ and thus, by using the analogue of [4, Proposition 3.2], which can be proved analogously as in [4] (again exploiting that, arguing as in [4, Lemma 2.17], the isoperimetric profile is achieved by bounded sets), we get
@@
On the other hand, for almost every radius, we have (see [6, Proposition 2.6], which can be applied thanks to Section˜2.3) that
@@
Conflicts of interest. The authors declare no conflicts of interest.